OpenStudy (anonymous):

What is the slope of the line given the equation 2x +7y=14 F. -2 G. - 7/2 H. - 2/7 J. 2

OpenStudy (texaschic101):

to find the slope of this line you would have to put the equation in y = mx + b form. Once you do this, you will find that the number in the m position is the slope. 2x + 7y = 14 -- subtract 2x from both sides 7y = -2x + 14 -- divide both sides by 7 y = -2/7x + 2 and as you can see, the number in the m position is -2/7, so that is your slope.

OpenStudy (anonymous):

Thankyou c:

OpenStudy (texaschic101):

no problem :)